OpenStudy (anonymous):

Solve for x. Round your answer to 2 decimal places.

OpenStudy (anonymous):



OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):



OpenStudy (anonymous):

help :/

OpenStudy (agent0smith):

Just solve that for x

OpenStudy (anonymous):

9.33

OpenStudy (agent0smith):

Not possible, the hyptenuse is always the longest side. cos51 = 12/x x = 12/cos51

OpenStudy (anonymous):

19.07